# Resolving Black Hole Family Issues Among the Massive Ancestors of Very High-Spin Gravitational-Wave Events Like GW231123

### v1

@@ Abstract
The latest detection of GW231123, a binary black hole merger with exceptionally large masses and high spins for the incoming components, has been suggested as a smoking gun for hierarchical formation. In this scenario, a first generation of black holes resulting from collapsing stars form in a dense environment such as a star cluster. Here they can assemble dynamically and undergo subsequent mergers. We discuss three challenges for the formation of a GW231123-like event inside a star cluster: (1) The high masses of the incoming black holes appear to be in the predicted pair-instability mass gap and thus suggest that second-generation or higher-order generation black holes are involved. (2) Very high spins () are very unlikely for dynamically assembled black holes because of the isotropic distribution of the spin vectors. (3) Hierarchically formed black holes are susceptible to receive large recoils, which could kick them out of their cluster and prohibit subsequent mergers. We simulate this scenario and show that only a few percent of mergers recover remnants within GW231123’s primary spin estimate (nearly all for the secondary) and are retained inside typical star clusters.
A large fraction of very rapidly spinning second-generation black holes (including ) can only form if the first-generation black holes merges with aligned spins. This is a natural outcome of massive binary star evolution scenarios, such as a chemically homogeneous evolution. This scenario also predicts equal masses for the components, implying that the resulting black holes tend to receive very low recoil kicks () and would therefore likely be retained inside a cluster. We conclude that GW231123-like events, if formed in a star cluster, could require first-generation black holes with large aligned spins that evolved through stellar binary interaction, followed by the dynamical assembly for a subsequent merger. We discuss the implications for the uncertain lower edge of the putative mass gap – and other alternative scenarios.
@@
Hierarchical mergers are expected in regions of large stellar densities such as star clusters [11, 12, 13, 14, 15, 16] or active galactic nuclei [17, 18, 19, 20], where gravitational few-body interactions facilitate the dynamical assembly and mergers of multiple generation of binary black holes. As such, hierarchical mergers have been invoked to explain the properties of a set of previously detected binary black hole mergers [17, 21, 13, 22, 23]. For GW231123 a hierarchical merger scenario in which both of its black holes were dynamically assembled is further supported by the inference of non-zero tilt angles between the black hole spins and the orbital angular momentum vector [1]. However, recovering the properties of the GW231123’s black holes through hierarchical mergers is not trivial considering the following challenges.
@@
Very large spin: Hierarchical mergers of dynamically assembled black holes yield remnant spin distributions centred around [26, 8]. This is consistent with the lower end of the uncertainty range of GW231123’s primary and well within the secondary’s range, but significantly below which is strongly supported by the LVK analysis [1].
@@
Cluster retention: Hierarchically formed black holes can receive large recoil kicks due to the asymmetric loss of linear momentum by gravitational-wave emission during the merger [27, 28, 29, 30]. The kick magnitude can be particularly large for unequal-mass mergers [31] and unequal spins [32] and may exceed the escape speed of a dense stellar environment [33, 34]. In particular, The LIGO Scientific Collaboration et al. [1, Sec. 6.5] argue GW231123’s properties would prefer the primary to be formed in a merger of black holes with masses and and the secondary from and . The recoils of such unequal masses would exceed the typical escape velocities of young star clusters and globular clusters [33] and would require GW231123 to occur in environments with significantly higher escape velocities.
@@ II Hierarchical mergers of black holes from binary star evolution
In the following, we consider a scenario in which very highly spinning massive black holes (such as GW231123’s primary) originate from the evolution of massive primordial binary stars in stellar clusters, which then, dynamically assemble and merge with another black hole in a GW231123-like event (see Figure 1). The exact binary fraction of (massive) binary stars in clusters is uncertain and depend on the cluster properties [e.g., 49]. Observational surveys show evidence that the binary abundance in globular clusters may be somewhat smaller than in the field of Galactic disk and halo, but not by a large factor [50]. While low-mass and wide binaries are likely to be disrupted by dynamical encounters, tight and massive binaries can remain bound [51]. Observations also exhibit a preference for equal-mass binaries and show that binaries are more commonly found in cluster centres than single stars, as a result of mass segregation [49, 52]. Recent observations suggest a higher fraction of binaries than predicted by simulations [52]; however, see also [53, 54] who reports lower fraction of short-period binaries than expected.
@@
We stress that the spin orientation of each black hole relative to the orbital angular momentum may change during the inspiral phase due to relativistic spin precession. As a result, the spin configuration established at large binary separations (for instance, when binary star evolution predicts nearly aligned spins) may not be preserved up to the moment of merger, which ultimately determines the remnant’s recoil and spin. However, for the equal-mass, aligned-spin systems we consider here, this deviation is expected to be negligible [see discussion in 43]. Furthermore, dynamical interactions with randomly passing stars in dense stellar environments may both accelerate the merger and torque the binary orbit away from perfect spin–orbit alignment. This effect can become significant in very dense environments, such as nuclear star clusters [55], especially when the delay time between binary black hole formation and merger is long [e.g., 55].
@@ III Discussion of binary star evolution scenarios
A natural formation channel for binary black holes with aligned spins is isolated binary evolution [see, e.g., 56, 57, 58]. Moreover, very massive black holes, which are the subject of interest for our scenario, are expected to undergo so-called direct collapse, which imparts little to no natal kick [59, 60, 58], leaving the spins aligned with the orbital angular momentum. Natal black hole spins predicted from isolated binary evolution are typically modest, with dimensionless spin parameters [56, 61] consistent as inferred so far by Abbott et al. [10]. This outcome suggests efficient angular momentum transport within stars [62, 63], in line with constraints on stellar rotation profiles derived from asteroseismic measurements [64, 65].
@@
The best observational candidates for highly spinning stellar-mass black holes come from X-ray binaries, where some black holes have estimated spin close to the critical value [e.g., 75, 76, 77]. However, their measured spin values may be systematically overestimated due to several factors; see the recent review by Zdziarski et al. [78].
We highlight that the presented scenario requires the formation of stellar-origin black holes with masses above M⊙ in order to reconcile the primary mass of GW231123, especially when accounting for the fact that a few percent of the total mass of the binary black holes are radiated away though gravitational waves. Such high masses were considered to lie within the so-called upper mass gap. In particular, pair-instability supernovae are expected to completely disrupt massive stars, leaving no remnant behind [3, 4, 5]. These processes were expected to produce a dearth of black holes within the mass range M⊙[72, 71, 79, 80]. Recent studies, however, have shown that the exact location of the mass gap depends sensitively on the highly uncertain 12C(,) 16O reaction rate, which can shift the edges of the gap to significantly lower or higher masses [24, 81, 25]. Adopting the lower limit within the 3-uncertainty of this reaction rate could potentially allow the formation of binary black hole mergers with component masses up to 90 – 100 M⊙ through isolated binary evolution [82, 83]. If the scenario presented here is realised in nature, it may have important implications for both the location of the upper mass gap and the underlying nuclear reaction rates. Alternatively, allowing for super-Eddington accretion during the second mass transfer phase could produce black hole masses above the pair-instability mass gap and result in a higher spin for one of the components [84, 85].
@@ IV Summary
The latest discovery of gravitational waves from the binary black hole merger GW231123 is exceptional not only because it involves the heaviest black hole masses and discovered so far, but also exhibits the strongest evidence for at least one very highly spinning component (). Large masses and spins are thought to be a smoking gun for a hierarchical formation scenario in which the black holes themselves were the outcome of previous mergers. However, reconciling the strong evidence for a very large spin of the primary is difficult through the canonical hierarchical merger scenario, where the primary would originate from dynamically assembled black holes with random spin directions. In this work, we explored various hierarchical formation scenarios to form the two black holes observed in GW231123. Assuming mutually random spin directions, we find that only – of all 1g+1g, 2g+1g, 2g+2g, and 3g+1g mergers, respectively, yield remnant black holes which are retained in a typical cluster environment () and remnant spins consistent with the lower limit of the measurement uncertainty of GW231123’s primary black hole.

### v2

@@ Abstract
The latest detection of GW231123, a black hole (BH) merger with exceptionally large component masses and high spins, has been suggested as a smoking gun for hierarchical formation. In this scenario, a first generation of BHs form from collapsing stars in dense environments such as star clusters, where they assemble dynamically and undergo subsequent mergers. We discuss three challenges for forming GW231123-like events in this scenario: (1) The high masses of the incoming BHs appear to be in the predicted pair-instability mass gap suggesting that higher-order generation BHs are involved. (2) Very high spins () are unlikely for dynamically assembled BHs because of the isotropic spin distribution. (3) Hierarchically formed BHs can receive large recoils that kick them out of their cluster and prohibit subsequent mergers. We simulate this scenario and show that only a few percent of mergers recover remnants within GW231123’s primary spin estimate and are retained inside typical star clusters.
A large fraction of very rapidly spinning second-generation BHs (including ) can form if the first-generation BHs merges with aligned spins. This is a natural outcome of massive binary star evolution scenarios, such as a chemically homogeneous evolution. This scenario also predicts equal masses for the components, implying that the resulting BHs receive low recoil kicks () and would therefore likely be retained inside a cluster. GW231123-like events, if formed in a star cluster, could require first-generation BHs with aligned spins formed from interacting stellar binaries, followed by the dynamical assembly for a subsequent merger.
@@
Hierarchical mergers are expected in regions of large stellar densities such as star clusters [87, 3, 71, 2, 30, 42, 67] or active galactic nuclei [103, 8, 100, 51], where gravitational few-body interactions facilitate the dynamical assembly and mergers of multiple generation of binary black holes. As such, hierarchical mergers have been invoked to explain the properties of a set of previously detected binary black hole mergers [103, 86, 71, 68, 102, 7, 66]. For GW231123 a hierarchical merger scenario in which both of its black holes were dynamically assembled is further supported by the inference of non-zero tilt angles between the black hole spins and the orbital angular momentum vector [95]. However, recovering the properties of the GW231123’s black holes through hierarchical mergers is not trivial considering the following challenges.
@@
Very large spin: Hierarchical mergers of dynamically assembled black holes are thought to yield remnant spin distributions centred around [40, 47]. This is consistent with the lower end of the uncertainty range of GW231123’s primary and well within the secondary’s range, but significantly below which is strongly supported by the LVK analysis [95].
@@
Cluster retention: Hierarchically formed black holes can receive large recoil kicks due to the asymmetric loss of linear momentum by gravitational-wave emission during the merger [25, 80, 18, 41]. The kick magnitude can be particularly large for unequal-mass mergers [52] and unequal spins [65] and may exceed the escape speed of a dense stellar environment [5, 48]. In particular, The LIGO Scientific Collaboration et al. [95, Sec. 6.5] argue GW231123’s properties would prefer the primary to be formed in a merger of black holes with masses and and the secondary from and . The recoils of such unequal masses would exceed the typical escape velocities of young star clusters and globular clusters [5] and would require GW231123 to occur in environments with significantly higher escape velocities, such as nuclear star clusters and active galactic nuclei, see e.g., Delfavero et al. [35].
@@ II Hierarchical mergers of black holes from binary star evolution
In the following, we consider a scenario in which very highly spinning massive black holes (such as GW231123’s primary) originate from the evolution of massive primordial binary stars in stellar clusters, which then, dynamically assemble and merge with another black hole in a GW231123-like event (see Figure 1). The exact binary fraction of (massive) binary stars in clusters is uncertain and depend on the cluster properties [e.g., 75]. Observational surveys show evidence that the binary abundance in globular clusters may be somewhat smaller than in the field of Galactic disk and halo, but not by a large factor [58]. While low-mass and wide binaries are likely to be disrupted by dynamical encounters, tight and massive binaries can remain bound [59]. Observations also exhibit a preference for equal-mass binaries and show that binaries are more commonly found in cluster centres than single stars, as a result of mass segregation [75, 50]. Recent observations suggest a higher fraction of binaries than predicted by simulations [50]; however, see also [76, 14] who reports lower fraction of short-period binaries than expected.
@@
The cumulative Gravitational-Wave Transient Catalog 4.0 [93] contains multiple merger events which lead to the formation of GW231123-like primary black holes and appear to be consistent with our formation scenario. Specifically, of the 218 merger events reported so far there are 18 whose final remnant black hole mass () measurements overlap with GW231123’s primary mass range. This number reduces to 13 if we only include events whose median final mass is which removes events with very large mass uncertainties. In Figure 4 we show their inferred posterior distributions of , , and . Most of these events have broad posterior distributions consistent with ( C.L.), none of them show unambiguous support for , and one event (GW231028 [93]) is strongly favouring a positive value . Recently, Antonini et al. [6] showed that the broadening of the distribution of black hole mergers with masses above 40 – is consistent with a subpopulation of mergers with large, isotropic spins from star clusters (see also Tong et al. [98] for a similar conclusion based on the sparsity of secondaries above those masses). We note that any massive merger with a clear support for negative detected in the future would further strongly support this scenario [84]. Here, we stress that the present sample which centres around is also consistent with some black hole spins being aligned and “both slowly spinning” or “slowly and highly spinning”, and may therefore consitute natural candidates to build up GW231123’s primary from primordial binary stars in a stellar cluster. In particular, GW231028’s support for positive strongly indicates that its black hole spins are co-aligned with the orbital angular momentum vector. Auspiciously, its final mass and final spin magnitude closely matches GW231123’s primary mass and spin, respectively, and we find () in 57 % (52 %) of its posterior samples [96], which is consistent with the “both highly spinning” model.
We stress that the spin orientation of each black hole relative to the orbital angular momentum may change during the inspiral phase due to relativistic spin precession. As a result, the spin configuration established at large binary separations (for instance, when binary star evolution predicts nearly aligned spins) may not be preserved up to the moment of merger, which ultimately determines the remnant’s recoil and spin. However, for the equal-mass, aligned-spin systems we consider here, this deviation is expected to be negligible [see discussion in 57]. Furthermore, dynamical interactions with randomly passing stars in dense stellar environments may both accelerate the merger and torque the binary orbit away from perfect spin–orbit alignment. This effect can become significant in very dense environments, such as nuclear star clusters [23], especially when the delay time between binary black hole formation and merger is long [e.g., 23].
@@ III Discussion of binary star evolution scenarios
A natural formation channel for binary black holes with aligned spins is isolated binary evolution [see, e.g., 16, 12, 10]. Moreover, very massive black holes, which are the subject of interest for our scenario, are expected to undergo so-called direct collapse, which imparts little to no natal kick [44, 61, 10], leaving the spins aligned with the orbital angular momentum. Natal black hole spins predicted from isolated binary evolution are typically modest, with dimensionless spin parameters [16, 20] consistent as inferred so far by Abbott et al. [1]. This outcome suggests efficient angular momentum transport within stars [91, 45], in line with constraints on stellar rotation profiles derived from asteroseismic measurements [29, 37].
@@
The best observational candidates for highly spinning stellar-mass black holes come from X-ray binaries, where some black holes have estimated spin close to the critical value [e.g., 74, 54, 55]. However, their measured spin values may be systematically overestimated due to several factors; see the recent review by Zdziarski et al. [105].
We highlight that the presented scenario requires the formation of stellar-origin black holes with masses above M⊙ in order to reconcile the primary mass of GW231123, especially when accounting for the fact that a few percent of the total mass of the binary black holes are radiated away though gravitational waves. Such high masses were considered to lie within the so-called upper mass gap. In particular, pair-instability supernovae are expected to completely disrupt massive stars, leaving no remnant behind [24, 56, 43]. These processes were expected to produce a dearth of black holes within the mass range M⊙[72, 70, 19, 90]. Recent studies, however, have shown that the exact location of the mass gap depends sensitively on the highly uncertain 12C(,) 16O reaction rate, which can shift the edges of the gap to significantly lower or higher masses [39, 31, 38]. Adopting the lower limit within the 3-uncertainty of this reaction rate could potentially allow the formation of binary black hole mergers with component masses up to 90 – 100 M⊙ through isolated binary evolution [21, 78]. If the scenario presented here is realised in nature, it may have important implications for both the location of the upper mass gap and the underlying nuclear reaction rates. Alternatively, allowing for super-Eddington accretion during the second mass transfer phase could produce black hole masses above the pair-instability mass gap and result in a higher spin for one of the components [101, 27]. In contrast, this scenario is challenged by recent works on the latest gravitational-wave data [93] by Antonini et al. [6] and Tong et al. [98] who interpret a broadening of for higher black hole masses and a dearth of massive secondary black holes, respectively, as the onset of a gap at – .
@@ IV Summary
The latest discovery of gravitational waves from the binary black hole merger GW231123 is exceptional not only because it involves the heaviest black hole masses and discovered so far, but also exhibits the strongest evidence for at least one very highly spinning component (). Large masses and spins are thought to be a smoking gun for a hierarchical formation scenario in which the black holes themselves were the outcome of previous mergers. However, reconciling the strong evidence for a very large spin of the primary is difficult through the canonical hierarchical merger scenario, where the primary would originate from dynamically assembled black holes with random spin directions. In this work, we explored various hierarchical formation scenarios to form the two black holes observed in GW231123. Assuming mutually random spin directions, we find that only – of all 1g+1g, 2g+1g, 2g+2g, and 3g+1g mergers, respectively, yield remnant black holes which are retained in a typical cluster environment () and remnant spins consistent with the lower limit of the measurement uncertainty of GW231123’s primary black hole. Auspiciously, we highlight recent work by Bamber et al. [11] who obtained very large remnants spins through collisional simulations of black hole clusters in full general relativity, suggesting it remnant spins to be more common. However, their specific choice of cluster properties and low-number statistics make it difficult to assess whether their cluster environment is a potential breeding ground for GW231123-like events.
@@
Later, several alternative scenarios have been proposed to explain the possible origin of GW231123. One class of models involves binary black hole mergers formed from isolated binaries of very massive stars above the upper pair-instability gap [32, 81, 53, 92]. In particular, high black hole spins and masses may arise from the chemically homogeneous evolution scenario [81]. [53] further demonstrate that massive low-metallicity stars with moderate magnetic fields can form black holes consistent in mass and spin with GW231123. In particular, in their models collapse above the pair-instability gap assisted by rotation and magnetic fields can drive mass loss through disk winds and jet launching, enabling the formation of highly spinning black holes within the gap.
Another alternative explanations invoke dynamical channels, however, under specific conditions and environments allowing for efficient accretion, such as active galactic nuclei disks [35, 13] or dense clusters [62]. Accretion from a companion or the surrounding medium in these environments can further increase black hole masses and spins [13, 62]. Nuclear star clusters can retain higher-generation black hole merger remnants despite recoil kicks, enabling multiple generations of black hole mergers. Delfavero et al. [35] found the observed masses and spins are most consistent with a merger between third- and fourth-generation black holes. Finally, some scenarios involving primordial black holes have also been suggested [104, 15, 33].